# PaliGemma: A versatile 3B VLM for transfer

### v1

@@ 3.2.2 Stage1: Multimodal pretraining
We train Stage1 at resolution 224px (hence, $N_{\text{img}}=256$ image tokens) and sequence length $N_{\text{txt}}=128$ for a total of 1 billion examples. While we provide an ablation in Section 5.1 showing that a 10x to 30x shorter Stage1 still provides good results on popular benchmarks, we wish to embue as much visual knowledge to the base model as possible, and cover a broad set of concepts, cultures, and languages [89, 82].
@@ Appendix A More related work
In the category of decoder-only VLMs, Fuyu [11] proposed to use an vision encoder-free architecture, that employs a transformer decoder-only model to process both image and text inputs. The input image is first patchfied and then linearly projected to a shared continuous embedding space as text tokens, so that the decoder-only model could process image and text tokens seamlessly. CM3 [2, 3, 125] and Chameleon [101] proposed to convert images into discrete tokens and then model them together with language in the token space with a shared transformer decoder model. Our work also shared promising early results of a Fuyu-style decoder-only setup following this line of work, by ablating the SigLIP vision encoder component from PaliGemma in Section 5.6.
@@ B.18 NLVR2: Reasoning about natural language grounded in photographs
B.19 OCR-VQA: VQA by reading text in images
@@
B.20 OKVQA: Outside knowledge VQA
@@
B.21 RefCOCO_seg: Referring expression segmentation
@@
B.22 RSVQA-hr: VQA for remote sensing (high res)
@@
B.23 RSVQA-lr: VQA for remote sensing (low res)
@@
B.24 SciCap: Captions for scientific figures
@@
B.25 ScienceQA: Science question answering
@@
B.26 Screen2Words: Mobile UI summarization
@@ B.27 ST-VQA: Scene text VQA
B.28 TallyQA: Complex counting questions
@@
B.29 CountBenchQA: Evaluate counting in a structured, controlled way
@@
B.30 TextCaps: Image captioning with reading comprehension
@@
B.31 TextVQA: Visual reasoning based on text in images
@@
B.32 VATEX-CAP: Broad video captioning
@@
B.33 VizWizVQA: VQA from people who are blind
@@
B.34 VQAV2: Visual Question Answering
@@
B.35 MMVP: Hard questions about CLIP-blind image pairs
@@
B.36 POPE: Object presence as yes/no VQA (object hallucination)
@@
B.37 Objaverse Multiview: view-consistent 3D object annotation
@@
B.38 WidgetCap: Descriptions for Mobile User Interface Elements
@@ Appendix C Image augmentations for RefCOCO
Figure 13 shows that one needs to explore both training-time resizing technique and inference-time strategy simultaneously. It seems natural that, when performing aspect-ratio preserving resize with padding at inference time, the training augemntations need to reflect that to some degree. However, one can also simply resize test images to a square at inference time, and then resize the output logits back to the original resolution for evaluation. Doing so works very well in tandem with simple resizing to square at training time and, as one might expect, does not work well with aspect-preserving resizes during training.
@@ Appendix I Inference
Table 5: Inference measurements on a TPUv3 [36] with 8 devices (4 chips). Even though FSDP sharding [132] is very efficient at training time (Section 3.2.6), at inference time it has the downside that each device needs process at least one single example and needs to read all model parameters. It reduces the memory requirements but it still processes one single example as slow as using a single device with a larger memory. Megatron-style sharding [95] on the other hand shards both the parameters and the activations, and can make use of multiple accelerators in parallel to process a single example and significantly reduce the total number of memory reads per device.
@@
FSDP float32 001 0130 014.7 03953 00068 28.8 00.58 03.48 01.69
008 0129 014.7 31656 00546 28.8 00.58 03.48 01.69
064 0782 014.7 41909 04365 38.2 04.46 02.72 01.87
512 6244 035.1 41984 14596 38.2 14.79 07.48 03.21
bfloat16 001 0112 008.3 04568 00121 33.3 01.01 01.75 00.64
008 0112 008.2 36516 00972 33.3 01.01 01.75 00.64
064 0779 010.2 42056 06249 38.3 06.36 01.87 00.81
512 6242 030.9 41994 16568 38.2 16.78 06.28 02.14
Megatron float32 001 0022 008.2 23014 00122 20.6 00.11 02.23 01.61
bfloat16 001 0017 005.5 30079 00181 26.9 00.27 00.92 00.82

### v2

@@ 3.2.2 Stage1: Multimodal pretraining
We train Stage1 at resolution 224px (hence, $N_{\text{img}}=256$ image tokens) and sequence length $N_{\text{txt}}=128$ for a total of 1 billion examples. While we provide an ablation in Section 5.1 showing that a 10x to 30x shorter Stage1 still provides good results on popular benchmarks, we wish to imbue as much visual knowledge to the base model as possible, and cover a broad set of concepts, cultures, and languages [68, 17, 93, 92, 85, 37, 136].
@@ Appendix A More related work
In the category of decoder-only VLMs, Fuyu [11] proposed to use an vision encoder-free architecture, that employs a transformer decoder-only model to process both image and text inputs. The input image is first patchified and then linearly projected to a shared continuous embedding space as text tokens, so that the decoder-only model could process image and text tokens seamlessly. CM3 [2, 3, 129] and Chameleon [105] proposed to convert images into discrete tokens and then model them together with language in the token space with a shared transformer decoder model. Our work also shared promising early results of a Fuyu-style decoder-only setup following this line of work, by ablating the SigLIP vision encoder component from PaliGemma in Section 5.6.
@@ B.18 NLVR2: Reasoning about natural language grounded in photographs
B.19 MaRVL: Reasoning about multilingual language grounded in multicultural photographs
[…]
Suffix: "True"
Train: Zero-shot evaluation of the model trained for NLVR2.
Metric: Accuracy (test splits).
Reference: Liu et al. [68]
B.20 OCR-VQA: VQA by reading text in images
@@
B.21 OKVQA: Outside knowledge VQA
@@
B.22 RefCOCO_seg: Referring expression segmentation
@@
B.23 RSVQA-hr: VQA for remote sensing (high res)
@@
B.24 RSVQA-lr: VQA for remote sensing (low res)
@@
B.25 SciCap: Captions for scientific figures
@@
B.26 ScienceQA: Science question answering
@@
B.27 Screen2Words: Mobile UI summarization
@@ B.28 ST-VQA: Scene text VQA
B.29 TallyQA: Complex counting questions
@@
B.30 CountBenchQA: Evaluate counting in a structured, controlled way
@@
B.31 TextCaps: Image captioning with reading comprehension
@@
B.32 TextVQA: Visual reasoning based on text in images
@@
B.33 VATEX-CAP: Broad video captioning
@@
B.34 VizWizVQA: VQA from people who are blind
@@
B.35 VQAV2: Visual Question Answering
@@
B.36 MMVP: Hard questions about CLIP-blind image pairs
@@
B.37 POPE: Object presence as yes/no VQA (object hallucination)
@@
B.38 Objaverse Multiview: view-consistent 3D object annotation
@@
B.39 WidgetCap: Descriptions for Mobile User Interface Elements
@@ Appendix C Image augmentations for RefCOCO
Figure 13 shows that one needs to explore both training-time resizing technique and inference-time strategy simultaneously. It seems natural that, when performing aspect-ratio preserving resize with padding at inference time, the training augmentations need to reflect that to some degree. However, one can also simply resize test images to a square at inference time, and then resize the output logits back to the original resolution for evaluation. Doing so works very well in tandem with simple resizing to square at training time and, as one might expect, does not work well with aspect-preserving resizes during training.
@@ Appendix I Inference
Table 5: Inference measurements on a TPUv3 [38] with 8 devices (4 chips), prefilling 512 tokens, cache sized 640 tokens. Even though FSDP sharding [137] is very efficient at training time (Section 3.2.6), at inference time it has the downside that each device needs process at least one single example and needs to read all model parameters. It reduces the memory requirements but it still processes one single example as slow as using a single device with a larger memory. Megatron-style sharding [99] on the other hand shards both the parameters and the activations, and can make use of multiple accelerators in parallel to process a single example and significantly reduce the total number of memory reads per device.
@@
FSDP float32 001 0107 014.6 04785 00068 34.9 00.6 02.04 01.69
008 0107 014.7 38282 00545 34.9 00.6 02.04 01.69
064 0772 014.7 42423 04367 38.6 04.5 02.67 01.87
512 6262 035.0 41866 14617 38.1 14.8 07.48 03.21
bfloat16 001 0101 008.2 05073 00122 37.0 01.0 00.84 00.64
008 0101 008.3 40555 00969 36.9 01.0 00.84 00.64
064 0764 010.2 42914 06258 39.1 06.4 01.48 00.81
512 6235 030.8 42043 16601 38.3 16.8 06.28 02.14
Megatron float32 001 0022 008.2 23018 00122 20.6 00.1 02.23 01.61
008 0105 010.3 39149 00777 34.9 00.7 02.32 02.01
064 0744 013.7 44055 04660 39.2 04.1 03.16 02.33
512 6360 041.2 41219 12427 36.7 11.5 07.82 03.75
bfloat16 001 0017 005.3 30006 00189 26.9 00.3 00.92 00.82
008 0099 006.5 41284 01239 36.7 01.2 01.00 00.80
064 0739 009.4 44348 06835 39.4 06.4 01.85 00.98
512 6348 034.9 41297 14677 36.7 13.1 06.54 02.41